[Hallc_running] [New Logentry] RC Daily Update

munoz at jlab.org munoz at jlab.org
Thu Dec 6 08:45:02 EST 2018

Logentry Text:
- Production data-taking continues with the SHMS at 9.79 degrees. 
- Last day was quite and productive. We took 2 h of dummy data at 40 uA, a couple of runs at 50 uA on LH2 and the rest at 25 uA.
- SHMS angle change planned for tomorrow morning.

RC Meeting:
- There will be an RC meeting today at 4:15pm, second floor counting house.

Notices in effect for shift crews:
- Shift leaders, please fill out the check list once per shift.
- Shift leaders, please do a quick cross-check of the BTA numbers against your shift summary.
- Check that the pre-scaler gui is active and hasn't frozen / doesn't need re-started.
- Be on the lookout for any errors when running the replay, particularly errors concerning missing reference times.
- Note the new beam positions for this energy (on the whiteboard / note on wall)
- NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''

Planned accelerator configuration changes:
- Energy change on Monday, 6am.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3635496
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181206/698bd43f/attachment.html>

More information about the Hallc_running mailing list